OpenStudy (anonymous):

9 over h ^{-3} Simplify?

OpenStudy (anonymous):

\[9 \over h ^{-3} \] Simplify?

OpenStudy (anonymous):

Remember: \[1/a^{-x} = a^x\] So can you go further or not??

OpenStudy (anonymous):

ummm? Let me see.

OpenStudy (anonymous):

I thought the expression was already simplified.

OpenStudy (anonymous):

you can take h in numerator by using the formula I have written above...

OpenStudy (anonymous):

i mean it look like this expression is already simplified to me.

OpenStudy (anonymous):

Ok that is really great..

OpenStudy (anonymous):

so it is?

OpenStudy (anonymous):

ugh O_o

OpenStudy (anonymous):

What you think what will be the answer???

OpenStudy (anonymous):

To expand upon what @waterineyes is stating: \[\frac{9}{h^{-3}}\] can be expressed as \[9*\frac{1}{h^{-3}}\] \[\frac{1}{h^{-3}}\] is an expression of \[\frac{1}{a^{-x}}\] and \[\frac{1}{a^{-x}} = a^{x}\]

OpenStudy (anonymous):

thsnk you!

OpenStudy (anonymous):

What's the answer you ended up with?

OpenStudy (anonymous):

9h^3

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

yay!! Thank you so much for taking the time to help me. ^_^